OpenStudy (anonymous):

How to find the vertex (x+4)(5x+2) so confused someone help me

OpenStudy (anonymous):

foil the binomial and consider the standard quadratic formula Ax^2+Bx+C=0 vertex= \[(-\frac{ b }{ 2a },f(-\frac{ b }{ 2a }))\]

OpenStudy (anonymous):

equation not formula

OpenStudy (johnweldon1993):

Right...so expand your 2 factors into a quadratic...and you'll take the 'b' value and divide it by 2 times the 'a' value... what is (x + 4)(5x + 2) when expanded?